# Identifying treatment response subgroups in observational time-to-event data

### v3

@@ B.4.4 Impact of treatment rate
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.790 (0.036) 0.036 (0.009) 0.034 (0.009) 0.018 (0.006) CSC Unadjusted 0.849 (0.033) 0.040 (0.005) 0.019 (0.005) 0.014 (0.004) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.149 (0.080) 0.188 (0.004) - 0.154 (0.005) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.808 (0.043) 0.035 (0.012) 0.027 (0.006) 0.019 (0.007) CSC Unadjusted 0.687 (0.048) 0.045 (0.011) 0.035 (0.016) 0.028 (0.006) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.010) 0.077 (0.005) 0.069 (0.004) CMHE ($L=3$) 0.318 (0.140) - - 0.132 (0.008) CMHE ($K=L$) 0.483 (0.104) 0.208 (0.013) - - KMeans + TE 0.002 (0.003) 0.195 (0.015) 0.079 (0.012) 0.148 (0.011) Virtual Twins 0.570 (0.013) 0.030 (0.005) 0.024 (0.013) 0.022 (0.005)
Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates.
B.4.5 Aligned covariates and treatment effect structure
@@
Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach. When covariates’ clusters are aligned with the outcome of interest, this method recovers the clustering structure well, as shown by the best Rand-Index. Non-randomisation of treatment, however, worsens the treatment estimates due to the absence of correction in the estimation of treatment responses.
@@
Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses. This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across settings.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.487 (0.052) 0.027 (0.013) 0.047 (0.020) 0.030 (0.009) CSC Unadjusted 0.522 (0.041) 0.025 (0.004) 0.027 (0.018) 0.034 (0.013) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.289 (0.078) 0.063 (0.009) 0.050 (0.019) 0.146 (0.020) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.666 (0.146) 0.016 (0.007) 0.029 (0.013) 0.033 (0.010) CSC Unadjusted 0.613 (0.207) 0.024 (0.016) 0.039 (0.015) 0.030 (0.003) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.294 (0.123) 0.057 (0.005) 0.061 (0.007) 0.144 (0.007) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
B.4.6 Linear treatment responses
As an alternative setting, we propose a setting with a linear treatment response, as assumed by CMHE Nagpal et al. (2023). To this end, we followed a similar process as described in Section B.1 at the exception of the treatment response. We draw the response under no treatment from a Gompertz, and the one under treatment with the same parameters shifted by the treatment effect. Formally, using the same notation
$B^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{0}_{z},A_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0},a_{z})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x)\times a_{z},s_{0}(\gamma_{z}^{0},x)\right)$
with $a_{z}$ the treatment effect for cluster $z$. In this experiments, we choose $a_{0},a_{1},a_{2}=(0.5,1,1.5)$.
Table 7 summarises performances in this scenario. While demonstrating a significant improvement of the CMHE alternatives, CSC presents competing performance on subgroup identification and treatment effect estimates.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.389 (0.065) 0.033 (0.015) 0.058 (0.035) 0.014 (0.006) CSC Unadjusted 0.411 (0.125) 0.029 (0.017) 0.062 (0.021) 0.017 (0.015) CMHE ($K=3$) 0.304 (0.033) 0.040 (0.003) 0.050 (0.003) 0.030 (0.003) CMHE ($L=3$) 0.211 (0.109) 0.051 (0.003) 0.106 (0.006) 0.028 (0.008) CMHE ($K=L$) 0.384 (0.060) 0.411 (0.797) - 0.435 (0.650) KMeans + TE 0.101 (0.139) 0.080 (0.011) 0.064 (0.028) 0.038 (0.023) Virtual Twins 0.265 (0.085) 0.025 (0.006) 0.018 (0.004) 0.064 (0.013) Observational CSC 0.311 (0.043) 0.029 (0.003) 0.074 (0.028) 0.027 (0.016) CSC Unadjusted 0.275 (0.168) 0.032 (0.005) 0.044 (0.044) 0.026 (0.014) CMHE ($K=3$) 0.360 (0.068) 0.038 (0.002) 0.050 (0.002) 0.035 (0.009) CMHE ($L=3$) 0.285 (0.111) 0.055 (0.006) 0.094 (0.007) 0.017 (0.009) CMHE ($K=L$) 0.382 (0.058) 0.052 (0.011) 0.128 (0.013) 0.028 (0.020) KMeans + TE 0.115 (0.154) 0.068 (0.020) 0.076 (0.031) 0.031 (0.008) Virtual Twins 0.186 (0.117) 0.048 (0.027) 0.041 (0.010) 0.039 (0.012)
Table 7: Cross-validated performance (with standard deviation in parenthesis) when considering a linear treatment response. This setting aligns with the assumption made by CMHE.
@@ B.4.6 Linear treatment responses
Appendix C Real-world case Study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
@@
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following Lee et al. (2018); Danks & Yau (2022); Jeanselme et al. (2023), we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  Danks & Yau (2022) for further description).
@@
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist Lazzari et al. (2023), making this setting more likely to meet the positivity assumption (assumption 6), necessary to study causality in observational data.
(a) Neural Survival Treatment
@@
Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning. As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
@@
Table 8 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) Royston & Parmar (2013). Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.00 (0.00) 93.3% 55.5% 1.18 (5.74) 17.4% 46.6% Subgroup 1 0.62 (0.16) 6.7% 47.5% 17.75 (16.08) 23.4% 48.6%

### v4

@@ B.5.3 Impact of treatment rate
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
@@
B.5.4 Aligned covariates and treatment effect structure
@@
Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach as this model makes this assumption. When covariates’ clusters are aligned with the outcome of interest, this method recovers the clustering structure well, as shown by the best Rand-Index.
@@
Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses. This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adverserial settings.
@@
Appendix C Case study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
@@
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following (Lee et al., 2018; Danks & Yau, 2022; Jeanselme et al., 2023), we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  (Danks & Yau, 2022) for further description).
@@
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist (Lazzari et al., 2023), making this setting more likely to meet the positivity assumption (Assumption 4), necessary to study causality in observational data.
(a) Causal Survival Clustering
@@
RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.01 (0.00) 94.6% 55.6% 1.2 (5.90) 17.5% 46.5% Subgroup 1 0.84 (0.11) 5.4% 45.1% 20.5 (15.56) 23.4% 50.4%
@@
Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE. For this latter, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold. As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
@@
Table 7 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) (Royston & Parmar, 2013). Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.